During an impact time lasting approximately 5.0*10^-4 s a golf club exerts an average force of about 500N on the golf ball. What linear impulse is delivered to the ball as a result?
1
Expert's answer
2020-10-23T11:50:59-0400

We can find the linear impulse delivered to the ball from the impulse-momentum change equation:


"F_{avg}\Delta t=m\Delta v,""F_{avg}\Delta t=mv_f-mv_i,""F_{avg}\Delta t=mv_f=p,"

here, "m" is the mass of the golf ball, "v_i=0" is the initial velocity of the golf ball, "v_f" is the final velocity of the golf ball, "F_{avg}=500\ N" is the average force that the golf club exerts on the golf ball, "\Delta t=5.0\cdot10^{-4}\ s" is the impact time during which the golf club interacts with the golf ball, "p" is the linear impulse delivered to the ball.

Then, from this equation we can find the linear impulse delivered to the ball:


"p=F_{avg}\Delta t=500\ N\cdot 5.0\cdot10^{-4}\ s=0.25\ \dfrac{kgm}{s}."

Answer:

"p=0.25\ \dfrac{kgm}{s}."